On a 200 km bike cycle ride, you ride the first 90 km at an average speed of 20m/s. What must your average speed over the next 40 km to be have your average speed for the total 200 km be?

a. 60m/s.

b. Given this average speed for the first 90 km, can you possibly attain an average speed of 80 m/s for the total 200 km ride? Explain.


1
Expert's answer
2021-01-30T06:25:35-0500

Answer

a) time taken for 90km (t)

="\frac{90000}{20}=45000" sec

Time taken for rest

(t') ="\frac{20000-90000}{x}=110000/x" sec

So average speed(given)

=total diatance/total time

"60=\frac{20000}{4500+\frac{110000}{x}}"

x="-\frac{110000}{1.167}m/s"

Speed is negative which is not possible.

b) now average speed 80m/s


"800=\frac{20000}{4500+\frac{110000}{x}}"

x=-55m/s


Speed is negative which is not possible.